please help me find the equation of the line {with slope -1/3 passing through (4, -1)}

Answer by Theo(13342)

slope intercept form of the equation of a straight line is:
y = mx + b
m is the slope
b is the y-intercept.
since the slope is -1/3, then the equation becomes:
y = -1/3 * x + b
the point is (4,-2)
that means that the x-coordinate is 4 and the y coordinate is -1.
to find b, replace y with -1 and x with 4 to get:
-1 = -1/3 * 4 + b
solve for b to get:
b = -1 + 4/3 = -3/3 + 4/3 = 1/3.
the equation becomes y = -1/3 * x + 1/3
here's the graph of that equation.
you can see that the line goes through the point (4,-1).

$$$
-